A student has money in three accounts that pay 5%, 7%, and 8% in annual interest. She has three times as much invested at 8% as she does at 5%. If the total amount she has invested is $2,700 and her interest for the year comes to $194, how much money does she have in each account?
Answer by mananth(16946)

A student has money in three accounts that pay 5%, 7%, and 8% in annual interest. She has three times as much invested at 8% as she does at 5%. If the total amount she has invested is $2,700 and her interest for the year comes to $194, how much money does she have in each account?

Let the student invest x amount in 5 %

5%------------7%-----------------8% --------- Total
x-------2700-(x+3x)--------------3x--------- 2700 Invested amount
The sum of all the 3 interests = 194
0.05x + 0.07* 2700-(x+3x)+ 0.08*3x = 194
0.05x +0.07*( 2700 -x-3x) +0.08*3x =194
0.05x + 189-0.28x + 0.24x=194
0.08x =194-189
x= 5/0.01
x= 500 in 5% account
3x = 1500 in 8 % account
2700-(500+1500) = 700 in 7 % account